OpenStudy (anonymous):

Suppose cosθ = 0.766 and 270° < θ < 360°. Find sinθ.

terenzreignz (terenzreignz):

okay... first, is sin θ going to be positive or negative in that interval?

OpenStudy (anonymous):

negative

terenzreignz (terenzreignz):

Okay, good. So we can use this identity \[\huge \cos^2\theta + \sin^2\theta = 1\]\[\huge \sin^2\theta = 1 - \cos^2 \theta\]

terenzreignz (terenzreignz):

plug in for cos θ

terenzreignz (terenzreignz):

and take the square root of both sides \[\huge \sin \theta\]

terenzreignz (terenzreignz):

\[\huge \sin \theta = -\sqrt{1-\cos^2 \theta}\] (remember that sine would be negative)

OpenStudy (anonymous):

so would \[\cos ^2\] be removed and replace with 0.766? or would it be 0.766^2?

terenzreignz (terenzreignz):

0.766^2 :) cos θ = 0.766 cos^2 θ = 0.766^2

OpenStudy (anonymous):

oooh ok

terenzreignz (terenzreignz):

Satisfied? :)

OpenStudy (anonymous):

yep. is the answer -0.643? just to be sure?

terenzreignz (terenzreignz):

Yes.

OpenStudy (anonymous):

ok. Thank you :)

terenzreignz (terenzreignz):

Good that you understood it :) I have to go now. -------------------------------------------- Terence out